OpenStudy (anonymous):

Can someone explain to me how to get the integral of 1/(x^2+2x+5)

OpenStudy (zarkon):

\[x^2+2x+5=(x+1)^2+4\]

OpenStudy (anonymous):

I understand that far, it is what follows where i get lost. I know i am supposed to end up with arc tan at the end too...

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

looks almost like \[\int \frac{dx}{x^2+a^2}=\frac{1}{a}\tan^{-1}(\frac{x}{a})\]

OpenStudy (misty1212):

cept you have \(x+1\) instead of \(x\) which really is not too important